# On the foundations and applications of Lorentz-Finsler Geometry

### v1

@@ 3.1.2. The Lorentz-Finsler setting
Starting at Fathi and Siconolfi’s article [55] published in 2012, a number of new approaches have been developed extending the existence of Cauchy time functions to cone structures $\mathcal{C}$, even lowering the regularity considered here. Indeed, they use cones just convex, closed, salient with non-empty interior and varying continuously and use weak KAM theory. Under global hyperbolicity, such a $\mathcal{C}$ is proved to admit a smooth Cauchy time function. Noticeably, there is a previous related approach in dynamical systems by Sullivan [157], as well as a posterior one by Monclair [130, 131], the latter linking attractors, chain recurrent points and the existence of time functions in Lorentzian manifolds by using Conley theory. This theory was used shortly after by Bernard and Suhr [28, 29] to prove the existence of temporal functions satisfying $h$-steepness (a condition which, in the Lorentz case, becomes stronger than Cauchy temporal but independent of steepness, see [34]). Their result holds for more general closed cone structures which requires a suitable extension of the notion of global hyperbolicity.
Summing up, at least for regular cone structures, the existence of Cauchy temporal functions as well as some of the properties listed above (as the stability in item (1) or the flexibility in item (2), see [124]), is well stablished now by methods alternative to the original Lorentzian ones. However, the possibility of splittings beyond the smooth one does not seem to have been explored. Next, we emphasize that the original approach [24, 25, 26, 134, 3] extends to Lorentz-Finsler manifolds yielding a global splitting and ensuring the properties listed above.
@@ 3.1.3. Case without boundary
Recall that, given a cone structure $\mathcal{C}$, a smooth function $\tau$ is temporal when $d\tau$ is timelike everywhere and then, the (pointwise) map $\sharp$ defines the (necessarily timelike) gradient $\nabla\tau=d\tau^{\sharp}$ (Remarks 2.10, 2.14).
@@
Any globally hyperbolic Finsler spacetimes admits a Cauchy temporal function $\tau$ and, then, a splitting $\mathds{R}\times S$ such that:
@@
  • $\nabla\tau$ is everywhere parallel to the natural vector field $\partial_{\tau}$.
Thus, the factors $\mathds{R}$ and $S$ of the product are everywhere $g_{\partial_{t}}$-orthogonal.
@@
The technique used in [25] to construct $\tau$ uses the following ingredients which can be transplanted to the Lorentz-Finsler setting [^14] Reference [147] is also strongly recommended for a proof in book format as well as further implications on the initial value problem in General Relativity..
First, Geroch’s Cauchy time function $t:M\rightarrow\mathds{R}$, whose existence in [68] is proved by using standard properties of causality valid for cone structures. Then, a smoothing procedure which relies on the existence of functions labelled $h_{p}$, $p\in M$ (introduced in [24]) defined as follows.
@@
The other steps in the references [24, 25] use algebraic manipulations of these functions that do not affect smoothability and, then, they are extensible to the Lorentz-Finsler case. It is worth pointing out that only finite sums of these types of functions are used in [24], while in [25] infinite sums are used. Anyway, no additional issue appears because the simple argument which proves that a converging sum of timelike vectors is also timelike (see [25, Lemma 3.10]) is valid for Lorentz-Minkowski norms too.
Once $\tau$ has been constructed, the splitting $\mathds{R}\times S$ is obtained by using that $\nabla\tau$ is a timelike vector field (this holds in the Lorentz-Finsler case by the consistency of Def. 2.9) and, then, its integral curves must cross all the slices of $\tau$. The last assertion follows because the splitting procedure flows $S=\tau^{-1}(0)$ by using the vector field $\nabla\tau/|\nabla\tau|^{2}$, then guaranteeing that $\partial_{t}$ and $\nabla\tau$ points out in the same oriented direction. ∎
@@ 3.1.4. Case with timelike boundary for $\mathcal{C}$ and summary
Now, all the notions like the definitions of inextensibility for causal curves $\gamma$, the sets $I^{+}(p),J^{+}(p)$ and the compactness of $J(p,q)$, must take into account the points of $\partial M$ and: (a) this forces to consider locally Lipschitz regularity instead of piecewise smoothness for the causal curves $\gamma$ used to compute $J^{\pm}(p)$ (by the reasons argued in [3, Appendix B], appliable even if $\partial M$ is $C^{\infty}$), (b) this permits to use $\partial M$ to extend $\gamma$ along it, and then (c) necessarily, Cauchy hypersurfaces $\bar{S}$ must have a boundary, eventually met by some inextensible $\gamma$ ’s.
@@
(c) Any Lorentz-Finsler metric $L$ endowed with a Cauchy temporal function $\tau$ adapted to the boundary admits a global Cauchy spliting $\bar{M}=\mathds{R}\times\bar{S}$ satisfying all the properties stated in Theorem 3.1.
@@
(a) Notice first that, in the Lorentzian case, the technique to obtain a (non-adapted) Cauchy temporal in the case without boundary also works for the case with boundary [3, Remark 4.3(2)]. Thus, this holds in the Lorentz-Finsler case by the reasons in Theorem 3.1.
(b) Choose any future directed timelike vector field $T$ on $\bar{M}$ such that $T|_{\partial M}$ is tangent to $\partial M$, [^15] This can be done by using that both $M$ and $\partial M$ are manifolds without boundary, thus admitting timelike vector fields $T_{M}$, $T_{\partial M}$, resp. for the inherited cone structures. $T_{\partial M}$ is tangent to $\partial M$ and it is also extendible to a timelike vector $T_{W}$ on neighborhood $W$ of $\partial M$. Then, $T=:\mu T_{M}+(1-\mu)T_{W}$, where $\{\mu,1-\mu\}$ is a partition of the unity subordinated to $\{M,W\}$, suffices. and assume $d\tau(T)=1$ with no loss of generality. Using Theorem 2.17, consider the associated cone triple $(\Omega=d\tau,T,F)$ and function $\mathcal{G}=d\tau^{2}-(F\circ\pi_{T})^{2}$.
By construction, at each point $p\in\bar{M}$, $T$ attains the minimum of $d\tau$ restricted to $\Sigma_{\mathcal{G}}|_{p}:=\mathcal{G}^{-1}(1)\cap{T_{p}M}$. Then, one can smoothen $\Sigma_{\mathcal{G}}$ (thus $\mathcal{G}$) close to the direction of $T$ to obtain a smooth and strongly convex $\Sigma$ following the procedure in [99, §5.2]. It is important to notice that this procedure permits to maintain $T$ as the direction of the minimum of $d\tau|_{\Sigma}$. Indeed, revising the procedure, it is enough to ensure that 0 remains as the minimum for the smoothed function $\tilde{t}_{0}$ in [99, Lemma 5.4]. This can be achieved directly by using the general argument by D. Azagra explained in [99, Remark 5.5]. Thus, taking the corresponding Lorentz-Finsler metric $L$ (with the so-smoothened indicatrix $\Sigma$), the gradient $\nabla\tau$ will point everywhere in the direction of $T$, in particular along $T(\partial M)$ on $\partial M$.
@@
The strenghetening of the conditions on temporal functions for the Lorentzian case in items (1) to (5) of §3.1.1 rely on arguments similar to those in the proof of Theorem 3.1. It is worth pointing out that cone triples $\Omega,T,F)$ provide a simple technical way to implement operations in cone structures as widening or narrowing cone strucuture (say, replacing $F$ by a Finsler metric with indicatrix enclosed or enclosed the one of $F$, resp.) to be used for stability. Thus, these five items can be extended to the Lorentz-Finsler case as before. Moreover, the strengthening in item (6) corresponds directly to the Lorentz-Finsler result in Theorem 3.3.
Notice that some of these stregthenings are relevant in its own, but items (3) and (4) are also important for applications to Nash-type isometric embeddings and convergence of Lorentz metrics. Thus, it would be interesting to explore Lorentz-Finsler issues in these fields, taking into account the subtleties of the purely Finsler problem (see for example [33] and [170]).
Finally, the analytic case in item (7) can also be extended to the Lorentz-Finsler case reasoning as in [150]. Namely, Grauert [74, Prop. 8] and Whitney [166, Lemmas 6, 7] (see also [5]) established that any $C^{k}$ function, $k\in\mathds{N}$, on a real analytic manifold can be $C^{k}$ approximated by analytic functions. Then, a $C^{2}$ approximation of a (eventually steep and $h$-steep) Cauchy temporal function by analytic ones suffices to maintain the spacelike character of the slices and its Cauchy character (apart from the other eventual properties), as well as to construct the orthogonal splitting.
@@ 3.2. The space of cone geodesics
The space $\mathcal{N}$ of null geodesics or lightrays or cone geodesics (the latter used here to emphasize its exclusive dependence on the cone structure), was introduced by Low in [113] for the Lorentz setting, see also his survey [114]. As we will see in §3.2.1, its basic framework can be extended directly to the Lorentz-Finsler setting, and it is natural to wonder at what extent the results in the Lorentz setting are also transplantable to the Lorentz-Finsler case. Moreover, as explained in §3.2.2, the case with timelike boundary ensures the good behavior of $\mathcal{N}$ for Finsler spacetimes admitting a timelike conformal boundary, in particular, conformally asymptotically anti-de Sitter spacetimes. It is worth pointing out the study in [79] of physical implications for $\mathds{N}$ to detect Finslerian elements.
@@ 3.2.1. Setting and Low’s conjecture
When the Finsler spacetime is globally hyperbolic, $\mathcal{N}$ becomes a smooth manifold. Indeed, choose any Cauchy hypersurface $S$, and notice that cone geodesics will cross $S$ exactly once, then providing an identification between $\mathcal{N}$ and the set of lightlike directions on $S$. This set can be characterized by taking any cone triple $(\Omega,T,F)$ for $\mathcal{C}$ and noticing that lightlike directions are in bijective corresponden with determined unit vectors for $F$, that is:
@@
When $\mathcal{C}$ is strongly causal, the quotient is a possibly non-Hausdorff manifold of the same dimension, see [113, Prop. 2.1] (or [84, Prop. 4.3]). Notice that non-Hausdorffness arises easily. Indeed, this happens in the space of (Riemannian) geodesics for $\mathds{R}^{2}\setminus\{0\}$ and this is transmited to the space of cone geodesics in $\mathds{R}^{1}_{1}\times(\mathds{R}^{2}\setminus\{0\})$ (an open subset of Lorentz-Minkowski ${\mathbb{L}}^{3}$). Anyway the structures to be considered next make sense whenever $\mathcal{C}$ is a (possibly non-Hausdorff) manifold.
As defined above, $\Sigma_{S}$ becomes also a convex bundle $\Sigma_{S}\rightarrow S$, however, a better insight for $\mathcal{N}$ appears regarding it as a quotient in the tangent bundle $T^{*}M$ as follows. In the Lorentz case, Low used the Legendre map (Def. 2.9) to first consider $\mathcal{C}^{\flat}$ as a subset of $T^{*}M$, second take quotients similar to those above and third regard $\mathcal{N}$ as a quotient of $\mathcal{C}^{\flat}$. Then, the standard contact form of $T^{*}M$ descend to $\mathcal{N}$ and it provides a primary tool linking the differential topology of $\mathcal{N}$ with the causal geometry of $M$ [114].
Indeed, the points of the manifold can be recovered as submanifolds in $\mathcal{N}$, namely, for each $x\in M$, define the sky $S_{x}\subset\mathcal{N}$ as:
$S_{x}=\{\gamma\in\mathcal{N}\mid x\in\gamma\},$
@@
Low conjectured that, in a globally hyperbolic spacetime, causal relations between events in $M$ can be detected by topological linking of their skies in $\mathcal{N}$. Subsequent work by Natário and Tod [136] clarified that the proper invariant for $n>2$ is Legendrian linking. Chernov and Nemirovski proved the conjecture in the following promiment case.
@@
For spacetimes with Cauchy hypersurfaces diffeomorphic to an open subset of $\mathds{R}^{n}$, $n>2$ (resp. $n=2$) and $q\in M, pp$ and $q$ are causally related if and only if $\mathfrak{S}_{p}$ and $\mathfrak{S}_{q}$ are nontrivially Legendrian linked (resp. topologically linked) in the contact manifold $\mathcal{N}$.
@@
An obvious natural question is to check whether the extension to the Lorentz-Finsler case holds. There are, however, other issues related to the topic which deserve attention as the role of null refocusing [104, 45, 14], causal boundary constructions [114, 15, 44], inverse and reconstruction problems [56, 13] or the space $\mathcal{N}$ in specific geometries [117].
3.2.2. The case with timelike boundary

### v2

@@ 3.1.2. The Lorentz-Finsler setting
Starting at Fathi and Siconolfi’s article [59] published in 2012, a number of new approaches have been developed extending the existence of Cauchy time functions to cone structures $\mathcal{C}$, even lowering the regularity considered here. Indeed, they consider cones just convex, closed, salient with non-empty interior and varying continuously, and they apply weak KAM theory. Under global hyperbolicity, such a $\mathcal{C}$ is proved to admit a smooth Cauchy time function. Noticeably, there is a previous related approach in dynamical systems by Sullivan [172], as well as a posterior one by Monclair [140, 141], the latter linking attractors, chain recurrent points and the existence of time functions in Lorentzian manifolds by using Conley theory. This theory was used shortly after by Bernard and Suhr [30, 31] to prove the existence of temporal functions satisfying $h$-steepness (as mentioned above, this condition becomes stronger than Cauchy temporal in the Lorentz case but independent of steepness, [36]). Their results hold for more general closed cone structures; with this purpose, a suitable extension of the notion of global hyperbolicity is introduced.
Summing up, at least for regular cone structures, the existence of Cauchy temporal functions and some of the properties listed in §3.1.1 (as the stability in item (1) or the flexibility in item (2), see [135]), is well stablished now by methods alternative to the original Lorentzian ones —however, the possibility of low regular splittings does not seem to have been explored. Next, we emphasize that the original approach [26, 27, 28, 144, 4] extends to Lorentz-Finsler manifolds yielding a global splitting and ensuring the properties listed in §3.1.1.
@@ 3.1.3. Case without boundary
Recall that, given a cone structure $\mathcal{C}$, a smooth function $\tau$ is temporal when $d\tau$ is timelike everywhere and then, the (pointwise) map $\sharp$ defines the (necessarily timelike) gradient $\nabla\tau=d\tau^{\sharp}$ (Remarks 2.11, 2.15).
@@
Any globally hyperbolic Finsler spacetime admits a Cauchy temporal function $\tau$ and, then, a splitting $\mathds{R}\times S$ such that:
@@
  • $\nabla\tau$ is everywhere collinear to the natural vector field $\partial_{\tau}$.
Thus, the factors $\mathds{R}$ and $S$ of the product are everywhere $g_{\partial_{t}}$-orthogonal; moreover, all its smooth Cauchy hypersurfaces are diffeomorphic.
@@
The technique used in [27] to construct $\tau$ uses the following ingredients which can be transplanted to the Lorentz-Finsler setting [^13] Reference [160] is also strongly recommended for a proof in book format as well as further implications on the initial value problem in General Relativity.. First, Geroch’s Cauchy time function $t:M\rightarrow\mathds{R}$, whose existence in [72] is proved by using standard properties of causality valid for cone structures. Second, a smoothing procedure which relies on the existence of functions labelled $h_{p}$, $p\in M$ (introduced in [26]) defined as follows.
@@
The other steps in the references [26, 27] use algebraic manipulations of these functions that do not affect smoothability and, then, they are extensible to the Lorentz-Finsler case. It is worth pointing out that only finite sums of these types of functions are used in [26], while in [27] infinite sums are used. Anyway, no additional issue appears because the simple argument which proves that a converging sum of timelike vectors is also timelike (see [27, Lemma 3.10]) is valid for Lorentz-Minkowski norms too.
Once $\tau$ has been constructed, the splitting $\mathds{R}\times S$ is obtained by using that $\nabla\tau$ is a timelike vector field (this holds in the Lorentz-Finsler case by the consistency of Def. 2.10) and, then, its integral curves must cross all the slices of $\tau$. The last assertion follows because the splitting procedure flows $S=\tau^{-1}(0)$ by using the vector field $\nabla\tau/|\nabla\tau|^{2}$, then guaranteeing that $\partial_{t}$ and $\nabla\tau$ points out in the same oriented direction; this also provides a diffeomorfism between any two smooth Cauchy hypersurfaces. ∎
@@ 3.1.4. Case with timelike boundary for $\mathcal{C}$ and summary
Now, all the notions like the definitions of inextensibility for causal curves $\gamma$, the sets $I^{+}(p),J^{+}(p)$ and the compactness of $J(p,q)$, must take into account the points of $\partial M$ and: (a) this forces to consider locally Lipschitz regularity instead of piecewise smoothness for the causal curves $\gamma$ used to compute $J^{\pm}(p)$ (by the reasons argued in [4, Appendix B], appliable even if $\partial M$ is $C^{\infty}$), and (b) necessarily, Cauchy hypersurfaces $\bar{S}$ must have a boundary, eventually met by inextensible $\gamma$ (which may travel along $\partial M$).
@@
(c) Any Lorentz-Finsler metric $L$ endowed with a Cauchy temporal function $\tau$ adapted to the boundary admits a global Cauchy splitting $\bar{M}=\mathds{R}\times\bar{S}$ satisfying all the properties stated in Theorem 3.1.
@@
(a) Notice first that, in the Lorentzian case, the technique to obtain a (non-adapted) Cauchy temporal function for the case without boundary also works for the case with boundary [4, Remark 4.3(2)]. Thus, this holds in the Lorentz-Finsler case by the reasons in Theorem 3.1.
(b) Choose any future directed timelike vector field $T$ on $\bar{M}$ such that $T|_{\partial M}$ is tangent to $\partial M$, [^14] This can be carried out just by noticing that both $M$ and $\partial M$ are manifolds without boundary, thus admitting timelike vector fields $T_{M}$, $T_{\partial M}$, resp. for the inherited cone structures. $T_{\partial M}$ is tangent to $\partial M$ and it is also extendible to a timelike vector $T_{W}$ on neighborhood $W$ of $\partial M$. Then, $T=:\mu T_{M}+(1-\mu)T_{W}$, where $\{\mu,1-\mu\}$ is a partition of the unity subordinated to $\{M,W\}$, suffices. and assume $d\tau(T)=1$ with no loss of generality. Using Theorem 2.18, consider the associated cone triple $(\Omega=d\tau,T,F)$ and function $\mathcal{G}=d\tau^{2}-(F\circ\pi_{T})^{2}$.
By construction, at each point $p\in\bar{M}$, $T$ attains the minimum of $d\tau$ restricted to $\Sigma_{\mathcal{G}}|_{p}:=\mathcal{G}^{-1}(1)\cap{T_{p}M}$. Then, one can smoothen $\Sigma_{\mathcal{G}}$ (thus $\mathcal{G}$) close to the direction of $T$ to obtain a smooth and strongly convex $\Sigma$ following the procedure in [107, §5.2]. It is important to notice that this procedure permits to maintain $T$ as the direction of the minimum of $d\tau|_{\Sigma}$. Indeed, revising the procedure, it is enough to ensure that 0 remains as the minimum for the smoothed function $\tilde{t}_{0}$ in [107, Lemma 5.4]. This can be achieved directly by using the general argument by D. Azagra explained in [107, Remark 5.5]. Thus, taking the corresponding Lorentz-Finsler metric $L$ (with the so-smoothened indicatrix $\Sigma$), the gradient $\nabla\tau$ will point everywhere in the direction of $T$, in particular along $T(\partial M)$ on $\partial M$.
@@
The strenghtening of the conditions on temporal functions for the Lorentzian case in items (1) to (5) of §3.1.1 rely on arguments similar to those in the proof of Theorem 3.1. It is worth pointing out that cone triples $(\Omega,T,F)$ provide a simple technical way to implement operations in cone structures as narrowing or widening them (say, replacing $F$ by a Finsler metric with indicatrix inside or enclosing the one of $F$, resp.) to be used for stability. Thus, these five items can be extended to the Lorentz-Finsler case as before. Moreover, the strengthening in item (6) corresponds directly to the Lorentz-Finsler result in Theorem 3.3.
Notice that some of these stregthenings are relevant in its own, but items (3) and (4) are also important for applications to Nash-type isometric embeddings and convergence of Lorentz metrics. Thus, it would be interesting to explore Lorentz-Finsler issues in these fields, taking into account the subtleties of the purely Finsler problem (see for example [35]).
Finally, the analytic case in item (7) can also be extended to the Lorentz-Finsler case reasoning as in [163]. Namely, Grauert [79, Prop. 8] and Whitney [184, Lemmas 6, 7] (see also [8]) established that any $C^{k}$ function, $k\in\mathds{N}$, on a real analytic manifold can be $C^{k}$ approximated by analytic functions. Then, a $C^{2}$ approximation of a (eventually steep and $h$-steep) Cauchy temporal function by analytic ones suffices to maintain the spacelike character of the slices and its Cauchy character (apart from the other eventual properties), as well as to construct the splitting in Th. 3.1.
@@ 3.2. The space of cone geodesics
The study of the space $\mathcal{N}$ of null geodesics or lightrays or cone geodesics (the latter used here to emphasize its exclusive dependence on the cone structure) comes from Penrose’s seminal ideas on twistor space, leading to consider lightrays as a fundamental physical structure from where spacetime emerges [152, 151]. Inspired by this approach, Low studied $\mathcal{N}$ introducing the viewpoint of contact geometry in the Lorentz setting [121, 123]. Hedicke’s thesis [85], developed the Lorentz-Finsler case providing background for this section.
Low proved that $\mathcal{N}$ turns out well behaved as a natural (Hausdorff) manifold in all globally hyperbolic spacetimes [121]. Hedicke [86] extended this result to some classes of causally simple spacetimes. The differentible structure of $\mathcal{N}$ can also be fully determined for globally hyperbolic cone structures-with-timelike-boundary, in particular, relativistic spacetimes admitting a timelike conformal boundary such as (conformally asymptotic) anti-de Sitter [91]. Low [123] also developed the symplectic and contact viewpoints connecting causality with topological linking, and posed a crucial conjecture solved in positive by Chernov and Nemirowski [46].
From a physical viewpoint, it is worth mentioning the study by Hasse and Perlick [84] about physical implications of $\mathcal{N}$ in order to detect possible Finslerian corrections to standard Relativity.
@@ 3.2.1. Setting and Low’s conjecture
When the Finsler spacetime is globally hyperbolic, $\mathcal{N}$ becomes a smooth manifold. Indeed, choose any Cauchy hypersurface $S$, and notice that cone geodesics will cross $S$ exactly once, then providing an identification between $\mathcal{N}$ and the set of lightlike directions on $S$. This set can be characterized by taking any cone triple $(\Omega,T,F)$ for $\mathcal{C}$ and noticing that lightlike directions are in bijective correspondence with the indicatrix for $F$ on $\Sigma$, that is, rewritting the conclusions in [121, 85]:
@@
When $\mathcal{C}$ is strongly causal, the quotient is a possibly non-Hausdorff manifold of the same dimension, see [121, Prop. 2.1] (or [91, Prop. 4.3]). Notice that non-Hausdorffness arises easily. Indeed, this happens in the space of (Riemannian) geodesics for $\mathds{R}^{2}\setminus\{0\}$ and this is transmited to the space of cone geodesics in $\mathds{R}^{1}_{1}\times(\mathds{R}^{2}\setminus\{0\})$ (an open subset of Lorentz-Minkowski ${\mathbb{L}}^{3}$). Anyway the structures to be considered next make sense whenever $\mathcal{C}$ is a (possibly non-Hausdorff) manifold.
Below strong causality, a class of cone structures with well behaved $\mathcal{N}$ is Zollfrei spacetimes [169], that is, compact Lorentzian manifolds such that the geodesic flow restricted to the cones induces a fibration by circles, thus, the orbit space being a smooth manifold.
As defined above, $\Sigma_{S}$ becomes also a convex bundle $\Sigma_{S}\rightarrow S$, the Hilbert contact structure (see §2.5.2) is preserved on $\mathcal{C}$ by the flows of $\mathcal{R},G$ and it can be induced on $\mathcal{N}$. Anyway, to provide a better insight, $\mathcal{N}$ will be regarded as a quotient in the cotangent bundle $T^{*}M$ by using the Legendre map (Def. 2.10). Following Low [123] in the Lorentz case (or [85, Sect. 3.1] in the Lorentz-Finsler one), first consider $\mathcal{C}^{*}$ as a subset of $T^{*}M\setminus\mathbf{0}$ (the $H_{L}$-energy level $E=0$), second take quotients similar to those above (which work specifically for $E=0$) and third regard $\mathcal{N}$ as a quotient of $\mathcal{C}^{*}$. Then, the standard contact form on $T^{*}M$ induces a contact form $\eta$ on $\mathcal{N}$ [85, Prop. 3.1.4]. This provides a tool which will link the differential topology of $\mathcal{N}$ with the causal geometry of $M$.
Indeed, the points of the manifold can be recovered as submanifolds in $\mathcal{N}$, namely, for each $p\in M$, define the sky $\mathfrak{S}_{p}\subset\mathcal{N}$ as:
$\mathfrak{S}_{p}=\{\gamma\in\mathcal{N}\mid p\in\gamma\},$
@@
Low conjectured that, in a globally hyperbolic spacetime, causal relations between events in $M$ can be detected by topological linking of their skies in $\mathcal{N}$. Subsequent work by Natário and Tod [146] clarified that the proper invariant for $n>2$ is Legendrian linking. Chernov and Nemirovski proved the reformulated conjecture in the following remarkable case.
@@
For spacetimes with Cauchy hypersurfaces diffeomorphic to an open subset of $\mathds{R}^{n}$, $n>2$ (resp. $n=2$) and $q\in M^{n+1}, pp$ and $q$ are causally related if and only if $\mathfrak{S}_{p}$ and $\mathfrak{S}_{q}$ are non-trivially Legendrian linked (resp. topologically linked) in the contact manifold $\mathcal{N}$.
@@
This result might attract attention from the Lorentz-Finsler setting, as well as other related topics, including the role of null refocusing [112, 48, 16], causal boundary constructions [123, 17, 47], inverse and reconstruction problems [60, 15] or the space $\mathcal{N}$ in specific geometries [126].
[…]
Moreover, Hedicke [87], inspired in ideas from symplectic geometry as Shelukhin’s Hofer norm on contactomorphisms [166], defined a Lorentzian type distance on the connected part of the identity $\hbox{Cont}_{0}(M,\eta)$. Among its nice properties, this distance is continuous, makes $\hbox{Cont}_{0}(M,\eta)$ a Lorentzian pre-length space and it can be calculated for the Reeb flow of the contact form —then permitting to answer a question by Shelukhin on the diameter of the contactomorphism group under the assumption of orderability [166, Question 18]. When applied to the space of skies of globally hyperbolic spacetimes, the manifold topology coincides with the topology induced by the distance of Legendrian skies in the space of cone geodesics [87, Th. 6.1].
3.2.2. Causal simplicity and the case with timelike boundary
The first natural question on $\mathcal{N}$ is its Haudorffness, only ensured above by global hyperbolicity. The step in the standard causal ladder of spacetimes just below global hyperbolicity is causal simplicity, that is, being causal with all $J^{+}(p),J^{-}(p)$ closed, [137]. Low characterized the compactness of the intersections $J^{+}(p)\cap J^{-}(p)$ for global hyperbolicity in terms of null pseudoconvexity [122], and Vatandoost et al. [179, Th. 2] gave a similar characterization of causal simplicity in terms of maximal null pseudoconvexity. Chernov [44] conjectured that causal simplicity might suffice for Hausdorfness. A partial positive answer was given by Hedicke and Suhr:
Theorem 3.8.
[88, Thm 2.5] If a causally simple spacetime can be isometrically embedded as an open subset of a globally hyperbolic spacetime then $\mathcal{N}$ inherits a Hausdorff contact manifold structure.
However, these authors also gave counterexamples to Chernov’s conjecture, see [88, §2] (a different one can be seen in [91, §5.2]). Indeed, they found examples of causally simple spacetimes non openly embedded in a globally hyperbolic spacetime with both Hausdorff and non-Hausdorff $\mathcal{N}$.
Anyway, causal simplicity turns out a natural class where the Hausdorffness of $\mathcal{N}$ should be studied. This class does not have a natural candidate for the manifold structure, in contrast to the globally hyperbolic case (Prop. 3.5). A natural subclass to focus is bounded convex open subsets of Lorentz-Minkowski space ${\mathbb{L}}^{n+1}$. Indeed, Hedicke [86] considered a more general class of open subsets for globally hyperbolic spacetimes, namely, the conformally star-shaped ones. For this class, $\mathcal{N}$ becomes contactomorphic to the space of cone geodesics of the surrounding globally hyperbolic spacetime, being then its contact type that of a spherical cotangent bundle (in particular, $ST^{*}\mathds{R}^{n}$ for the case of ${\mathbb{L}}^{n+1}$), see [86, Th. 2.3].
Another interesting class of open neighborhoods, the strongly null convex ones, has also been considered in [86]. The results are restricted to dimension 3, where they connect to other parts of contact geometry. For such a neighborhood $K$, the boundary $\partial\mathcal{N}_{K}$ of its space of cone geodesics $\mathcal{N}_{K}$ becomes a convex surface included in the cone geodesic space $\mathcal{N}$ of the globally hyperbolic spacetime [86, Th. 2.8]. Here, convex surface is used in the contact sense introduced by Giroux [78], namely, a closed surface $S$ with a transverse contact vector field defined on a neighborhood of $S$ (this author also proved that these surfaces are generic among closed surfaces), thus, the two notions of convexity are linked. Under further hypotheses, $\mathcal{N}_{K}$ becomes contactomorphic to the spherical cotangent bundle of $\mathds{R}^{2}$, [86, Th. 2.9].